OpenStudy (anonymous):

The cost of 5 shirts and 3 pants is $105.00. The cost of 2 shirts is $9.00 more than the cost of one pair of pants. What is the cost of each? I need to solve this using elimination.

OpenStudy (anonymous):

Do you hv difficulty in forming the equations?? Can you solve the equations on your own??

OpenStudy (anonymous):

Yes I have difficulties with equations. I don't know why, but I always have.

OpenStudy (anonymous):

OK... Let the cost of 1 shirt = $x and cost of 1 pant = $ y Based on this can you form two equations...........

OpenStudy (anonymous):

5x+3y=105? I'm not sure how to get the other equation either. I know it's x+y equals something.

OpenStudy (anonymous):

The cost of 5 shirts and 3 pants is $105.00. The cost of 2 shirts is $9.00 more than the cost of one pair of pants. What is the cost of each? Let the cost of 1 shirt = $x and cost of 1 pant = $ y The cost of 5 shirts and 3 pants is $105.00. --- this gives us 5x + 3y = 105 The cost of 2 shirts is $9.00 more than the cost of one pair of pants. ---- this gives us 2x = 1y + 9 this can be rearranged to get second equation as 2x - 1y = 9

OpenStudy (anonymous):

now that you hv the two eqns, solve them to get your values of x and y...

OpenStudy (anonymous):

x=12 and y=15

OpenStudy (anonymous):

Can you proceed further on your own???

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

it was the equation part i was having trouble with

OpenStudy (anonymous):

Good!! Your ans is correct...........

OpenStudy (anonymous):

Ok